# Belief Dynamics Reveal the Dual Nature of In-Context Learning and Activation Steering

### v1

@@ 4.2 Altering model belief: Effects of Activation Steering
$p(c_{i}|x)=p(c_{i}|v)=\sigma\left(-w^{\intercal}v-b\right)=\sigma\bigg(-\beta_{i}\|d_{i}\|^{2}-\sum_{j\neq i}\beta_{j}d_{i}^{T}d_{j}-b\bigg)\approx\sigma\left(-\beta_{i}(v)a-b\right),$ (6)
where $a=\|d_{i}\|^{2}$. Using this, we can express the posterior odds as follows: $\log\frac{p(c_{i}|v)}{p(c_{i}^{\prime}|v)}=\log\frac{p(c_{i}|v)}{1-p(c_{i}|v)}=a\beta_{i}(v)+b.$ Thus, if one steers the model representation along direction $d_{i}$, e.g., changing $v$ to $v+m\cdot d_{i}$ [^2] Note that steering boundlessly (e.g., taking $m\to\infty$) will push the representations to a region that lies outside the support over which distribution $P(c|v)$ is defined. We see such effects empirically (e.g., see Fig. 12) and thus focus our discussion in the main paper in a range for $m$ where posterior-belief changes monotonically., the model’s belief in concept $c_{i}$ will linearly increase (in log-space) to $a\beta_{i}(v)+a\cdot m+b$. Relating this back to the unsteered model’s log-posterior odds, we get the following:
$\log\frac{p(c_{i}|v+m\cdot d_{i})}{p(c_{i}^{\prime}|v+m\cdot d_{i})}=\log\frac{p(v|c_{i})}{p(v|c_{i}^{\prime})}+\log\frac{p(c_{i})}{p(c_{i}^{\prime})}+a\cdot m=\log\frac{p(v|c_{i})}{p(v|c_{i}^{\prime})}+\log\frac{p^{\prime}(c_{i})}{p^{\prime}(c_{i}^{\prime})}\ \ .$ (7)
That is, steering yields a constant shift in the model log-posterior odds that will consistently change model beliefs for both an individual observation $x$ or an entire population $X\sim P_{x}$. We therefore argue the effects of steering are best described as alteration of a model’s prior beliefs in a concept $c$, updating the log prior odds from $\log\frac{p(c)}{p(c^{\prime})}$ to $\log\frac{p^{\prime}(c)}{p^{\prime}(c^{\prime})}$ (where $p^{\prime}(c)$ is an unnormalized prior). Intuitively, this formalizes the claim that steering vectors should be expected to change behavior $y$ regardless of the input $x$. For example, for the concept $C_{\text{happy}}$ we should expect the steering vector $\hat{d}_{c}$ to make an LM behave more happy even if it is given inputs $x^{(c^{\prime})}$ that are not happy, i.e., which have lower $p(c\ |\ x^{(c^{\prime})})$.
@@
Prediction 2  Assuming linear representation hypothesis holds, Eq. 7 shows steering will increase a model’s belief in concept $c_{i}$ at a sigmoidal rate with steering magnitude $m$.
Figure 7: Predicting the amount of context needed to enable a persona. (Top) The belief dynamics model is highly predictive of crossover points $N^{*}$ between $c$ and $c^{\prime}$, with a correlation of $r=0.97$, and (Bottom) our $N^{*}$ estimates effectively predict phase boundaries in empirical behavioral data.
@@ 4.3 Final model
From Eq. 7 , the log posterior odds given an intervention on $v$ can be defined as $\log o(c|x)=\log\frac{p(c)}{p(c^{\prime})}+\log\frac{p(x|c)}{p(x|c^{\prime})}+a\cdot m$, where $o(v|c)=o(x|c)$ since we assume $p(c|x)=p(c|v)$ in Sec. 4.2. Next, we substitute the log prior odds $\log\frac{p(c)}{p(c^{\prime})}$ with a constant offset $b$, since it does not depend on the precise input $x$ or its representation $v$. This gives us our final model of belief update dynamics in ICL:
@@
This model describes how model behavior changes as a function of both context length $N$ and steering magnitude $m$. Concretely, for the model prediction results described in this work, we fit scalar parameters to $a$, $b$, $\gamma$, $\alpha$ to empirical averages of model behavior $p(y|x)=p(c|x)$ (Eq. 4) using L-BFGS, across various contexts $x$ (where $N=|x|$) and steering with various magnitudes $m$. We do so for practical reasons: first, the steering vectors we use in practice are not the true concept vectors $d_{i}$, and so the effect of steering magnitude will be $a\propto\|d_{i}\|^{2}$ (but not necessarily $a=\|d_{i}\|^{2}$), and second, we estimate the prior odds $b$ rather than observing the concept priors $p(c),p(c^{\prime})$. This model allows us to compute the transition points in context length for a given steering magnitude $m$ when the model’s belief in concept $c$ surpasses $c^{\prime}$, i.e., when $\log o(c|x)=0$:
@@ Results
Observing the phase diagrams in Fig. 6, we find that our model is highly predictive of the joint effects of in-context learning and steering. Moreover, following our definition of $N^{*}(m)$ in Eq. 9, we can predict the crossover points when behavior will transition to be dominated by $c$ (Fig. 7).
@@ 5 Discussion
We see there being a number of exciting directions for future follow-up work. The findings in our work may have practical implications for model control, to help practitioners understand how to best combine behavioral methods like ICL with mechanistic interventions for the purpose of controlling language models. The phase boundaries we find across ICL and steering could also have important consequences for AI safety, since language model behavior might suddenly and dramatically change after some threshold of context or steering is passed. Predicting these transition points, as we have done in this work, may prove to be an essential tool for safe and effective control of language models. One limitation of this work is that we only consider binary concepts and use only one method for constructing steering vectors (CAA). Future work may explore how our theory and model generalize to non-binary concept spaces, where there may be more than one direction for belief to vary across. Another compelling direction for future work is to explore how belief is updated with alternative steering vector methods such as SAEs (Templeton et al., 2024). In order to better understand the intersection of ICL and activation steering, another experiment we performed was to compute steering vectors over multiple shots (see App. D). We found that, counter-intuitively, after vectors were normalized to equal magnitude, steering vectors computed over multiple shots had an even weaker effect than steering with vectors computed over a single query. Further work is required to explain this effect. Finally, we also hope to explore in future work how activation steering and ICL interact with larger and more capable LLMs. Overall, by revealing a common Bayesian mechanism linking prompting and activation steering, our work re-frames how we understand belief and representation in LLMs, opening a rich space for theoretical exploration and principled model control in future work.
@@ A.1 Derivation of the Bayes Factor
    Concept log-likelihood declines proportionally to the number of mismatched labels: The persona-adoption settings we examine consist of query-label examples where labels are binary and either map or do not map to a persona. Thus, it is reasonable that the log-likelihood for a concept will decline proportionally with the number of mismatched labels seen. Assuming this likelihood function follows Goodman et al. [2008], who studied rule-based concept learning in humans. Formally, we can express the likelihood function for a concept $c$ as:
@@
    Log-likelihood scales as a power-law with number of in-context examples $N$: This assumption aims at accommodating the power-law behavior observed in studies of LLM in-context learning [Anil et al., 2024, Liu et al., 2024b]. Specifically, we assume the common form of a scaling-law from scaling laws predicting loss during pretraining $L(n)\approx L(\infty)+\frac{A}{n^{\alpha}}$ [Kaplan et al., 2020]. However, in our case, $L(N)$ represents the negative-log likelihood for the $N$-th in-context example [Anil et al., 2024]. Given this assumption, we derive a sub-linear discount term $\tau$ that arises from the ratio between the negative log-likelihood for $N$ in-context examples under the power-law assumption, and the negative log-likelihood given by an optimal Bayesian agent using the likelihood function from assumption 1. Following the derivation from Wurgaft et al. [2025], we write:
@@
A.2 Derivation of the Effect of Steering Magnitude (Eq. 7)
Here we show how the log posterior odds (Eq. 7) can be represented as:
@@ Appendix C Full Steering Range
The results discussed in our main text focus on the case where the Linear Representation Hypothesis (LRH) holds. However, we empirically find that with larger enough steering magnitudes, the linear effect of steering on $\log o(c|x)$ begins to break down and the sigmoidal response function we show in Fig. 5 converges towards 0 (Fig. 12 and Fig. 13). This is similar to the findings of Panickssery et al. [2024] which shows that LLM behavior begins to break down and become incoherent with very large magnitude steering vectors. We find that behavior converges towards chance ($p(y|x)=0.5$), even with very large context lengths.
@@
Appendix D Many-shot Steering Vector Computation
@@
Figure 14: Computing Steering Vectors Over Varying Number of Shots. Steering vectors are computed for Llama-3.1-8B for the Psychopathy dataset over varying number of shots. Note that here 0-shot refers to providing the model with a single target query and a "Yes/No" reply and taking the difference in mean activations, whereas a larger number of shots refers to the number of in-context examples provided to the model before the target query. Top panel shows the effect of steering vectors computed over varying number of shots and applied at different magnitudes and context lengths. Bottom panel shows cosine similarity between vectors computed over varying number of shots with the 0-shot vector or the 128-shot vector.
Appendix E Experimental Details
@@
In our experiments, we use Llama-3.1-8B-Instruct, Gemma-2-9B-Instruct, Qwen-2.5-7B. For efficiency reasons, we restrict our analyses to LLMs which balance relatively small scale ($\sim$ 8 billion parameters) with relatively high performance on major benchmarks. We use 4-bit quantization for further efficiency, and run inference locally, primarily on A100 GPUs. Steering vector training and application are implemented using an open-source repository  [^3] https://github.com/steering-vectors/steering-vectors for LLM steering, which implements Contrastive Activation Addition [Turner et al., 2024].

### v2

@@ 4.2 Altering model belief: Effects of Activation Steering
$p(c_{i}|x)$ $=p(c_{i}|v)=\sigma\left(-w^{\intercal}v-b\right)$ (7)
$=\sigma\Big(-\beta_{i}\|d_{i}\|^{2}-\sum_{j\neq i}\beta_{j}d_{i}^{T}d_{j}-b\Big)$
$\approx\sigma\left(-\beta_{i}(v)a-b\right),$
where $a=\|d_{i}\|^{2}$. Using this, we can express the posterior odds as follows: $\log\frac{p(c_{i}|v)}{p(c_{i}^{\prime}|v)}=\log\frac{p(c_{i}|v)}{1-p(c_{i}|v)}=a\beta_{i}(v)+b.$ Thus, if one steers the model representation along direction $d_{i}$, e.g., changing $v$ to $v+m\cdot d_{i}$ [^1] Steering boundlessly (e.g., taking $m\to\infty$) will push the representations to a region that lies outside the support over which distribution $P(c|v)$ is defined. We see such effects empirically (e.g., see Fig. 13) and thus focus our discussion in the main paper in a range for $m$ where posterior-belief changes monotonically., the model’s belief in concept $c_{i}$ will linearly increase (in log-space) to $a\beta_{i}(v)+a\cdot m+b$. Relating this back to the unsteered model’s log-posterior odds, we get the following:
$\log\frac{p(c_{i}|v+m\cdot d_{i})}{p(c_{i}^{\prime}|v+m\cdot d_{i})}$ $=\log\frac{p(v|c_{i})}{p(v|c_{i}^{\prime})}+\log\frac{p(c_{i})}{p(c_{i}^{\prime})}+a\cdot m$
$=\log\frac{p(v|c_{i})}{p(v|c_{i}^{\prime})}+\log\frac{p^{\prime}(c_{i})}{p^{\prime}(c_{i}^{\prime})}\ \ .$ (8)
That is, steering yields a constant shift in the model log-posterior odds that will consistently change model beliefs for both an individual observation $x$ or an entire population $X\sim P_{x}$. We argue the effects of steering are best described as alteration of a model’s prior beliefs in a concept $c$ by updating the log prior odds from $\log\frac{p(c)}{p(c^{\prime})}$ to $\log\frac{p^{\prime}(c)}{p^{\prime}(c^{\prime})}$ (where $p^{\prime}(c)$ is an unnormalized prior). Intuitively, this formalizes the claim that steering vectors should change behavior $y$ regardless of the input $x$. For example, for the concept $C_{\text{happy}}$ we should expect the steering vector $\hat{d}_{c}$ to make an LM behave more happy even with inputs $x^{(c^{\prime})}$ that are not happy, i.e., which have lower $p(c\ |\ x^{(c^{\prime})})$.
@@
Prediction 2  Assuming linear representation hypothesis holds, Eq. 4.2 shows steering will increase a model’s belief in concept $c_{i}$ at a sigmoidal rate with steering magnitude $m$.
Figure 7: Additional tasks. We find similar results for high prior concepts (App. E) and flipped-label sentiment analysis (App. F)
@@ 4.3 Final model
From Eq. 4.2 , the log posterior odds given an intervention on $v$ can be defined as $\log o(c|x)=\log\frac{p(c)}{p(c^{\prime})}+\log\frac{p(x|c)}{p(x|c^{\prime})}+a\cdot m$, where $o(v|c)=o(x|c)$ since we assume $p(c|x)=p(c|v)$ in Sec. 4.2. Next, we substitute the log prior odds $\log\frac{p(c)}{p(c^{\prime})}$ with a constant offset $b$, since it does not depend on the precise input $x$ or its representation $v$. This gives us our final model of belief update dynamics in ICL:
@@
This model describes how model behavior changes as a function of both context length $N$ and steering magnitude $m$. Concretely, for the model prediction results described in this work, we fit scalar parameters to $a$, $b$, $\gamma$, $\alpha$ to empirical averages of model behavior $p(y|x)=p(c|x)$ (Eq. 4.1) using L-BFGS, across various contexts $x$ (where $N=|x|$) and steering with various magnitudes $m$. We do so for practical reasons: first, the steering vectors we use in practice are not the true concept vectors $d_{i}$, and so the effect of steering magnitude will be $a\propto\|d_{i}\|^{2}$ (but not necessarily $a=\|d_{i}\|^{2}$), and second, we estimate the prior odds $b$ rather than observing the concept priors $p(c),p(c^{\prime})$.
Figure 8: Predicting the amount of context needed to enable a persona. (Top) The belief dynamics model is highly predictive of crossover points $N^{*}$ between $c$ and $c^{\prime}$, with a correlation of $r=0.97$ ($p<.001$), and (Bottom) our $N^{*}$ estimates effectively predict phase boundaries in empirical behavioral data.
This model allows us to compute the transition points in context length for a given steering magnitude $m$ when the model’s belief in concept $c$ surpasses $c^{\prime}$, i.e., when $\log o(c|x)=0$:
@@ Results
Observing the phase diagrams in Figs. 6,  7, we find that our model is highly predictive of the joint effects of in-context learning and steering. Moreover, following our definition of $N^{*}(m)$ in Eq. 10, we can predict the points when behavior will transition to be dominated by $c$ (Fig. 8).
@@ 5 Discussion
We see there being a number of exciting directions for future follow-up work. The findings in our work may have practical implications for model control, to help practitioners understand how to best combine behavioral methods like ICL with mechanistic interventions for the purpose of controlling language models. The phase boundaries we find across ICL and steering could also have important consequences for AI safety, since language model behavior might suddenly and dramatically change after some threshold of context or steering is passed. Predicting these transition points, as we have done in this work, may prove to be an essential tool for safe and effective control of language models. One limitation of this work is that we only consider binary concepts and use only one method for constructing steering vectors (CAA). Future work may explore how our theory and model generalize to non-binary concept spaces, where there may be more than one direction for belief to vary across. Another compelling direction for future work is to explore how belief is updated with alternative steering vector methods such as SAEs (Templeton et al., 2024). In order to better understand the intersection of ICL and activation steering, another experiment we performed was to compute steering vectors over multiple shots (see App. G). We found that, counter-intuitively, after vectors were normalized to equal magnitude, steering vectors computed over multiple shots had an even weaker effect than steering with vectors computed over a single query. Further work is required to explain this effect. Finally, we also hope to explore in future work how activation steering and ICL interact with larger and more capable LLMs. Overall, by revealing a common Bayesian mechanism linking prompting and activation steering, our work re-frames how we understand belief and representation in LLMs, opening a rich space for theoretical exploration and principled model control in future work.
@@ A.1 Derivation of the Bayes Factor
    Concept log-likelihood declines proportionally to the number of mismatched labels: The persona-adoption settings we examine consist of query-label examples where labels are binary and either map or do not map to a persona. Thus, it is reasonable that the log-likelihood for a concept will decline proportionally with the number of mismatched labels seen. Assuming this likelihood function follows Goodman et al. (2008), who studied rule-based concept learning in humans. Formally, we can express the likelihood function for a concept $c$ as:
@@
    Log-likelihood scales as a power-law with number of in-context examples $N$: This assumption aims at accommodating the power-law behavior observed in studies of LLM in-context learning (Anil et al., 2024; Liu et al., 2024a). Specifically, we assume the common form of a scaling-law from scaling laws predicting loss during pretraining $L(n)\approx L(\infty)+\frac{A}{n^{\alpha}}$ (Kaplan et al., 2020). However, in our case, $L(N)$ represents the negative-log likelihood for the $N$-th in-context example (Anil et al., 2024). Given this assumption, we derive a sub-linear discount term $\tau$ that arises from the ratio between the negative log-likelihood for $N$ in-context examples under the power-law assumption, and the negative log-likelihood given by an optimal Bayesian agent using the likelihood function from assumption 1. Following the derivation from Wurgaft et al. (2025), we write:
@@
A.2 Derivation of the Posterior Approximation (Eq. 7)
Here, we show how the posterior $p(c_{i}|x)$ for a concept $c_{i}$ given data $x$ can be approximated as a function of $\beta_{i}(v)$, which measures the degree to which concept vector $d_{i}$ is present in a given vector $v$: $p(c_{i}|x)\approx\sigma\left(-\beta_{i}(v)a-b\right)$.
First, since $v$ is a hidden representation corresponding to input $x$, we have:
$p(c_{i}|x)=p(c_{i}|v)$
Next, since LRH argues that for a linearly represented concept, a logistic classifier $\sigma\left(-w^{\intercal}v-b\right)$ can infer the extent to which $c$ is present in $v$:
$p(c_{i}|v)=\sigma\left(-w^{\intercal}v-b\right)$
For our third step, recall that by the LRH (Eq. 6), a given hidden representation $v$ can be decomposed into a sum of vectors $d_{i}$ for each concept $i$, weighted by the extent to which each concept is present in $v$: $v=\sum_{i}\beta_{i}(v)d_{i}$ (note that we abbreviate $\beta_{i}(v)$ here as $\beta_{i}$). This gives us:
$-w^{\intercal}v-b$ $=\sum_{j}\beta_{j}\ d_{i}^{\intercal}d_{j}-b$
[…]
$\approx-\beta_{i}(v)a-b$
Thus, by substituting this value into $\sigma\left(-w^{\intercal}v-b\right)$, we have:
$p(c_{i}|x)\approx\sigma\left(-\beta_{i}(v)a-b\right)$
A.3 Derivation of the Effect of Steering Magnitude (Eq. 4.2)
Here we show how the log posterior odds (Eq. 4.2) can be represented as:
@@ Appendix C Full Steering Range
The results discussed in our main text focus on the case where the Linear Representation Hypothesis (LRH) holds. However, we empirically find that with larger enough steering magnitudes, the linear effect of steering on $\log o(c|x)$ begins to break down and the sigmoidal response function we show in Fig. 5 converges towards 0 (Fig. 13 and Fig. 14). This is similar to the findings of Panickssery et al. (2024) which shows that LLM behavior begins to break down and become incoherent with very large magnitude steering vectors. We find that behavior converges towards chance ($p(y|x)=0.5$), even with very large context lengths.
@@
Appendix D Results for a Larger LLM
We tested whether our model can account for the in-context and steering dynamics of a larger LLM, Llama-3.1-70B. We find that, as with smaller LLMs, in-context learning and steering jointly affect behavior, and that we are able to predict behavior across varying context lengths and steering magnitudes with a high correlation (cross-validated out-of-sample $r=0.98$, see Fig. 15). In contrast with smaller models, ICL occurs substantially faster, and the models begin to match the persona with merely a few examples, whereas for smaller models in many cases more examples are required.
Figure 15: In-context learning and steering jointly affect behavior in Llama-3.1-70B. Note that, due to our cross-validation procedure operating over rows, as in previous figures, we plot the steering response function for Bayesian models trained over the full dataset, while the heatmap and correlation coefficients reflect cross-validated out-of-sample predictions.
Appendix E High-prior concepts
In addition to low-prior personas such as Psychopathy, we wanted to test whether our model can account for positive persona traits such as openness, which should in principle show similar dynamics only with a higher prior. We indeed find that we can capture these dynamics with a high correlation (cross-validated out-of-sample $r=0.87$, see right panel of Fig. 16), and that, despite the high-prior for persona consistent answer, the LLM’s behavior remains steerable, and the steering response function qualitatively retains the shape it does for lower-prior concepts (see Fig. 16), which is expected by our theory.
Figure 16: Our Bayesian Model Captures the Dynamics of High-Prior Concepts. Note that, due to our cross-validation procedure operating over rows, as in previous figures, we plot the steering response function for Bayesian models trained over the full dataset, while the heatmap and correlation coefficients reflect cross-validated out-of-sample predictions.
Appendix F Flipped-Label Sentiment Detection
In addition to persona settings, we also tested our model in a flipped-label sentiment detection setting, used by (Agarwal et al., 2024), in which a LLM learns a flipped mapping between sentences and sentiment labels over a financial sentiment analysis dataset (Malo et al., 2014a), as shown in the top left panel of Fig. 17. In this setting, unlike in the persona settings, there are three response labels. The steering is done between the standard label space and the flipped label space, and we measure "Label flipping behavior" as the probability assigned to the flipped label. We find that even in a setting containing non-binary labels, our model is able to predict LLM behavior across varying steering magnitudes and in-context shots with high accuracy (cross-validated out-of-sample $r=0.96$, see top right panel of Fig. 17). In contrast with the persona settings, the LLM requires substantially more shots in order to attain the label-flipping behavior.
Figure 17: Our Bayesian Model Captures LLM behavior in a Flipped-Label Sentiment Detection Setting. The top left panel illustrates the setting, in which LLMs view sentences and have to match them to sentiment labels, which are corrupted according to the illustrated rotation. We compute steering vectors using a difference-in-means between the rotated and standard label space over the sentiment analysis dataset. Due to our cross-validation procedure operating over rows, as in previous figures, we plot the steering response function for Bayesian models trained over the full dataset, while the heatmap, ICL curves and correlation coefficients reported reflect cross-validated out-of-sample predictions.
Appendix G Many-shot Steering Vector Computation
@@
Figure 18: Computing Steering Vectors Over Varying Number of Shots. Steering vectors are computed for Llama-3.1-8B for the Psychopathy dataset over varying number of shots. Note that “0-shot” here refers to providing the model with a single target query and a "Yes/No" reply and taking the difference in mean activations, whereas a larger number of shots refers to the number of in-context examples provided to the model before the target query. Top panel shows the effect of steering vectors computed over varying number of shots and applied at different magnitudes and context lengths. Bottom panel shows cosine similarity between vectors computed over varying number of shots with the 0-shot vector or the 128-shot vector.
Appendix H Experimental Details
@@
In our experiments, we use Llama-3.1-8B-Instruct, Gemma-2-9B-Instruct, Qwen-2.5-7B-Instruct, and Llama-3.1-70B-Instruct (elsewhere, “-Instruct” is omitted for brevity). For efficiency reasons, we primarily analyze LLMs which balance relatively small scale ($\sim$ 8 billion parameters) with relatively high performance on major benchmarks. We use 4-bit quantization for further efficiency, and run inference locally, primarily on A100 GPUs. Steering vector training and application are implemented using an open-source repository  [^2] https://github.com/steering-vectors/steering-vectors for LLM steering, which implements Contrastive Activation Addition (Turner et al., 2024).